OpenStudy (superfly123):

Write the equation of a line that is: Perpendicular to y=3/2x and whose y intercept is 5 parallel to y=2x+7 and which passes through (-2,5)

OpenStudy (superfly123):

For the first one i got y=-2/3x+5 The second one: y=2x+9

jigglypuff314 (jigglypuff314):

Perfect! :)

OpenStudy (superfly123):

But, the last equation when i plug in two points it doesn't make sense

OpenStudy (superfly123):

it doesnt match up with 2x + 7

jigglypuff314 (jigglypuff314):

if you plug in 5 for y and -2 for x into your answer equation 5 = 2(-2) + 9 5 = -4 + 9 5 = 5 --> true so it works ^_^

jigglypuff314 (jigglypuff314):

they don't care about the 2x+7 one, that was just there for you to get the slope :)

OpenStudy (superfly123):

Oh, so not necessarily they need to match up with the original one

jigglypuff314 (jigglypuff314):

mhmm ^_^

OpenStudy (superfly123):

Thanks